OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. B = 81°, b = 26, c = 28

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

we can see by trying the law of sines

OpenStudy (anonymous):

\[\frac{\sin(C)}{28}=\frac{\sin(81)}{26}\] so \[\sin(C)=\frac{28\sin(81)}{26}\]

OpenStudy (anonymous):

if \(\frac{28\sin(81)}{26}>1\) then there is no triangle because sine is always less than or equal to one

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=28sin%2881%29%2F26 as you can see, it is too large so no triangle

OpenStudy (anonymous):

yeah it's larger than 1 so it's no triangle right!?

OpenStudy (anonymous):

@satellite73